calculate the pH of a solution formed by mixing 40ml of 0.1M HCL with 60 ml of water.
1
Expert's answer
2011-12-02T08:47:22-0500
m(HCl)=0,1mol\L*36,5g\mol*0,04L=0,146g;
C(HCl)=0,146g : (36,5g\mol*(0,04L+0,06L))=0,04M=0,04mol\L;
For water [H+] = (Kw)1/2 = 10-7 mol\L.
Then, pH = - lg[H+] = - lg(10-7 +4*10-2) = 1,4.
pH = 1,4.